OpenStudy (anonymous):

Solve –4x + 7 = x + 2

OpenStudy (anonymous):

-4x -x = 2-7 -5x = -5 x=1

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

anytime ;)

mathslover (mathslover):

First of all : \[\huge{\mathbb{WELCOME}\textbf{TO}\mathbb{OPEN}\textbf{STUDY}}\]

OpenStudy (anonymous):

ngaww what a cute banner @mathslover yes welcome indeed!

mathslover (mathslover):

:D .. thanks @virtus